Question 1 Oriole Company has two divisions; Sporting Goods and Sports Gear. The sales mix is 65% for Sporting Goods and 35% for Sports Gear. Oriole incurs $6752500 in fixed costs. The contribution margin ratio for Sporting Goods is 30%, while for Sports Gear it is 50%. The break-even point in dollars is o $18250000. $2498425. O $16881250 $15703489. Click if you would like to Show Work for this question:
Answer #1

Weighted Contribution margin=Respective Contribution margin*Respective sales mix

=(0.65*0.3)+(0.35*0.5)=0.37

Hence breakeven point=fixed cost/weighted Contribution margin

=(6752500/0.37)

=$18,250,000.